OpenStudy (anonymous):

Two quadratic functions are shown. Function 1: f(x) = 2x2 − 8x + 1 Function 2: x g(x) −2 2 −1 −3 0 2 1 17 Which function has the least minimum value and what are its coordinates? Function 1 has the least minimum value and its coordinates are (0, 1). Function 1 has the least minimum value and its coordinates are (2, −7). Function 2 has the least minimum value and its coordinates are (0, 2). Function 2 has the least minimum value and its coordinates are (−1, −3).

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

any help

OpenStudy (campbell_st):

ok.... in the 2nd function, g(x) what is the lowest value in the g(x) column... and what x value is associated with it

OpenStudy (anonymous):

ok is it -2,2 @campbell_st

OpenStudy (campbell_st):

no... look down the g(x) column 1st... what is the lowest value...?

OpenStudy (anonymous):

-3

OpenStudy (campbell_st):

great and the x value associated with is it -1 so g(x) has a minimum at (-1, -3)

OpenStudy (campbell_st):

so is this algebra or calculus..?

OpenStudy (anonymous):

thank uu

OpenStudy (anonymous):

algebra

OpenStudy (campbell_st):

well you have to look at f(x) as well

OpenStudy (campbell_st):

ok for f(x) you need to find the line of symmetry, the minimum value lies on the line of symmetry so use \[x = \frac{-b}{2 \times a}\] your f(x) has b = -8 and a = 2 can you substitute them and calculate a value for x

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

not quite its x = 2 \[x = \frac{-(-8)}{2 \times 2} = 2\] so substitute x = 2 into f(x) and calculate a value... this will be the minimum for f(x)

OpenStudy (anonymous):

im confused

OpenStudy (campbell_st):

find \[f(2) = 2\times 2^2 - 8 \times 2 + 1\]

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok... so now which point is lower (2, -7) or (-1, -3)...?

OpenStudy (anonymous):

2,-7

OpenStudy (campbell_st):

there's your answer

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

u think u can help with a few more

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

no, i'm about to go to school... its friday morning in Australia

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thx anyways